A $1,000 corporate bond has a maturity date 20 years from now and a coupon rate of 6 percent paid annually. Calculate the value of the bond if the required rate of return is a) 4% b) 6% c) 8%.

Strip the bond into an interest only bond and a face value only bond. That is, create a bond that consists only of the coupon interest payments and one that consists only of the face value. Calculate the value of each of these new bonds with required rates of return of a) 4% b) 6% c) 8%. Which bond has more interest rate risk? Which bond has more default risk? Now say that instead of annual coupon payments, the corporation makes payments every six months. Recalculate the value of your stripped bonds. Explain the results.
Answer #1

Value of the bond is present value of all future cash outflows in the form of interest

Given information Amount in ($)

Corporate bond 1000 , Maturity Period 20 Years , Coupon rate is 6 % that mean interest is 60

Value of bond if Required rate of return is 4 %

= Interest * Cumulative Present value factor of 4% for 20 years

= 60 * 13.59

= 815.40

Value of bond if Required rate of return is 6 %

= Interest * Cumulative Present value factor of 6% for 20 years

= 60 * 11.47

= 688.20

Value of bond if Required rate of return is 8 %

= Interest * Cumulative Present value factor of 8% for 20 years

= 60 * 9.82

= 589.20